a parabola has its focus (-6,4) and x=-1 as directrix. Find its equation
Answer by Bxbblezone (1)

The vertex of the parabola is (-3.5, 4).
The focal length, p=-2.5 The parabola opens to the left.
(y-4)² = 4p(x+3.5)
(y-4)² = -10(x+3.5)
(y-4)² = -10x - 35